Helium is used to inflate football-shaped balloons that will be released in Memorial stadium for the opening game of the football season. One 2.00 liter balloon pressurized to 2.00 atm will float nicely on a hot day (35.0◦C). Unfortunately, a cold front moves through on opening day, and the temperature dips to 15.0◦C! The balloons look like PICKLES instead of footballs. What is the new volume of each balloon?
Answer in units of liters.
1
Expert's answer
2020-03-02T06:55:10-0500

pV=nRT n=pV/RT = 2 x 2 / 0.082 x 308 = 0.1583 moles.

V= 0.1583 x 0.082 x 288 / 2 = 1.87 L.